# Dionysos.jl: a Modular Platform for Smart Symbolic Control

### v1

@@ 1 Introduction
To address this issue, we introduce Dionysos.jl [^1] The package is open source and available on GitHub at: https://github.com/dionysos-dev/Dionysos.jl., a modular Julia [9] package for solving optimal control problems using state-of-the-art control techniques, from control, optimization, and machine learning, for complex systems. It is built on top of different Julia packages such as JuMP.jl and MathOptInterface.jl, and features optimal control problem definitions and several abstraction-based methods to solve them. Dionysos.jl is the software of the ERC project Learning to control (L2C) [^2] European Reasearch Council (ERC) under the European Union’s Horizon 2020 research and innovation program under grant agreement No 864017 - L2C.. It implements new fundamental techniques that have been designed for smart abstraction [16, 30, 20, 6, 15] as part of the L2C project, with the goal of designing control techniques that would come with guarantees of safety and efficiency, while at the same time being able to take into account non-standard constraints or information, e.g. coming from first principles in physics, a logical specification translating some legal regulations, or some recommendation from a human being.
@@ 2.1 Control framework
A tuple $(\boldsymbol{x},\boldsymbol{u})\in\mathcal{X}^{T+1}\times\mathcal{U}^{T}$ is a trajectory of length $T$ of the system $\mathcal{S}=(\mathcal{X},\mathcal{U},F)$ starting at $x(0)$ if $T\in\mathbb{N}\cup\{\infty\}$, $x(0)\in\mathcal{X}$, $\forall k\in\{0,\ldots,T\}:u(k)\in\mathcal{U}(x(k))$ and $x(k+1)\in F(x(k),u(k))$. The set of trajectories of $\mathcal{S}$ is called the behavior of $\mathcal{S}$, denoted $\mathcal{B}(\mathcal{S})$.
@@
We now define the control problem. {defi} Consider a system $\mathcal{S}=(\mathcal{X},\mathcal{U},F)$. A specification $\Sigma$ for $\mathcal{S}$ is defined as any subset $\Sigma\subseteq(\mathcal{X}\times\mathcal{U})^{\infty}$. It is said that system $\mathcal{S}$ satisfies the specification $\Sigma$ if $\mathcal{B}(\mathcal{S})\subseteq\Sigma$. A system $\mathcal{S}$ together with a specification $\Sigma$ constitute a control problem $(\mathcal{S},\Sigma)$. Additionally, a controller $\mathcal{C}$ is said to solve the control problem $(\mathcal{S},\Sigma)$ if $\mathcal{C}\times\mathcal{S}$ satisfies the specification $\Sigma$.
@@ 2.3 Smart abstraction
Figure 3: Piecewise constant controller $\forall x_{1}\in R^{-1}(x_{2}):\ \mathcal{C}_{1}(x_{1})=\{u\}$ where $u\in\mathcal{U}_{2}\subseteq\mathcal{U}_{1}$. Left: The two-dimensional concrete system with its state space discretization. Right: The abstract system.
Figure 4: Piecewise affine controller $\forall x_{1}\in R^{-1}(x_{2}):\ \mathcal{C}_{1}(x_{1})=\{\kappa(x_{1})\}$ where $\kappa\in\mathcal{U}_{2}$ and $\kappa(x_{1})=K_{\kappa}x_{1}+\ell_{\kappa}\in\mathcal{U}_{1}$. The map $G_{1}(x)=F_{1}(x,\kappa(x))$ provides the closed-loop transition under the controller $\kappa$. Left: The two-dimensional concrete system with its state space discretization. Right: The abstract system.
@@
which requires that all states in the initial set $\mathcal{X}_{I}$ remain (safely) in the set $\mathcal{X}_{S}$ forever. In addition to the specified specifications, a state cost function $c:\mathcal{X}\rightarrow\mathbb{R}_{+}$ evaluating the cost of being in a state $x$, and a transition cost function $t:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}_{+}$ quantifying the cost of transitioning from one state $x$ to another with control input $u$, can be supplied to the control problem. The objective is then to design a controller that satisfies the specification while minimizing the cumulative cost. : Discretization templates: The quantizer $R$ can either be a (partial or not) partition or cover of the concrete state space. In addition, given a continuous state space $\mathcal{X}\subseteq\mathbb{R}^{n}$, Dionysos.jl supports two types of sets for the quantizer:
  • a hyperrectangle of center $c\in\mathbb{R}^{n}$ and half-lengths $h\in\mathbb{R}^{n}_{+}$: $\operatorname{H}(c,h)=\{x\in\mathbb{R}^{n}\mid|x_{i}-c_{i}|\leq h_{i}\text{ for }i=1,\ldots,n\};$
  • an ellipsoid with center $c\in\mathbb{R}^{n}$ and shape defined by $P\in\mathbb{S}_{+}^{n}$ is denoted as $\operatorname{E}(c,P)=\{x\in\mathbb{R}^{n}\mid(x-c)^{\top}P(x-c)\leq 1\}.$
@@ 4.1 The System module
The System module contains mathematical descriptions of controlled systems (see Definition 2.1), controllers (see Definition 2.1) and trajectories. It is an extension of MathematicalSystems.jl and HybridSystems.jl [29], it complements the latter with more specific system definitions.
@@ 4.2 The Problem module
Note that, in Dionysos.jl, the obstacles are encoded as part of the domain of the system. The second structure is SafetyProblem{S,XI,XS,T<:Real}, and implements an invariance problem. Its fields are
@@ 4.3 The Optim module
A simple example of an implementation of a classical abstraction-based method is given in the following. In Code LABEL:lst:optimizer_optimize!, the steps given in Figure 1 are followed. In Code LABEL:lst:optimizer_struct, the structure for the optimizer is defined. To be initialized, it needs a concrete problem to solve, as well as a discretization for the state space and the input space. {lstlisting}[ language = Julia, numbers=left, label=lst:optimizer_struct, caption=Definition of an abstraction-based strategy structure implementing the abstract type MOI.AbstractOptimizer. ] mutable struct OptimizerT<:MOI.AbstractOptimizer concrete_problem::Union Nothing, PR.OptimalControlProblem, PR.SafetyProblem abstract_problem::Union Nothing, PR.OptimalControlProblem, PR.SafetyProblem abstract_controller::Any concrete_controller::Any state_grid::Any input_grid::Any function OptimizerT(cp,sg,ig) where T return newT( cp, nothing, nothing, nothing, nothing, sg, ig ) end end {lstlisting}[ language = Julia, numbers=left, label=lst:optimizer_optimize!, caption=Definition of the MOI.optimize! function for the strategy Optimizer defined in Code LABEL:lst:optimizer_struct. ] function MOI.optimize!(opt::Optimizer) # Build the abstraction abstract_system = build_abs_system( opt.concrete_problem.system, opt.state_grid, opt.input_grid, ) # Build the abstract problem abstract_problem = build_abs_problem( opt.concrete_problem, abstract_system ) opt.abstract_problem = abstract_problem # Solve the abstract problem abstract_controller = solve_abs_problem( abstract_problem ) opt.abstract_controller = abstract_controller # Solve the concrete problem opt.concrete_controller = solve_conc_problem( abstract_system, abstract_controller ) end
@@ 5 Numerical example
Let $E(c,P)\subseteq\mathbb{R}^{n}$ be an ellipsoid with center $c\in\mathbb{R}^{n}$ and shaped defined by $P\in\mathbb{R}^{n\times n}$, that is
$E(c,P):=\{x\in\mathbb{R}^{n}:(x-c)^{\top}P(x-c)\leq 1\}.$ (8)
@@
The system in Example 5, as well as other examples, are already defined in Dionysos.jl/problems. The OptimalControlProblem defining the problem in Example 5 can be found in the NonLinear module in problems/non_linear.jl. In Code LABEL:lst:ex_system, we import this problem. {lstlisting}[ language = Julia, numbers=left, label=lst:ex_system, caption=The problem stated above is imported from the problems directory. ] concrete_problem = NonLinear.problem() concrete_system = concrete_problem.system
@@
We choose to use AB.LazyEllispoidsAbstraction (see Table 1) to solve this problem, a smart abstraction method that constructs a memoryless concretization relation ($\operatorname{MCR}$) that partially covers the state space with ellipsoids. In Code LABEL:lst:ex_solver, we instantiate the corresponding Optimizer, then set all the needed fields, including concrete_problem, the reach-avoid problem stated above. For the sake of conciseness, we do not state the purpose of the other fields, and we gather them in the variable other_parameters in the code. {lstlisting}[ language = Julia, numbers=left, label=lst:ex_solver, caption=The Optimizer is instantiated, and the solver parameters are set. ] optimizer = MOI.instantiate( AB.LazyEllipsoidsAbstraction.Optimizer ) AB.LazyEllipsoidsAbstraction.set_optimizer!( optimizer, concrete_problem, other_parameters… )
@@
In Code LABEL:lst:ex_solve, we solve the problem and retrieve the corresponding abstract system, abstract problem and concrete controller (see Figure 1). Note that every operation follows the MathOptInterface.jl syntax. {lstlisting}[ language = Julia, numbers=left, label=lst:ex_solve, caption=The problem is solved, and the result is extracted from the solver. ] MOI.optimize!(optimizer) abstract_system = MOI.get( optimizer, MOI.RawOptimizerAttribute("abstract_system") ) abstract_problem = MOI.get( optimizer, MOI.RawOptimizerAttribute("abstract_problem") ) abstract_controller = MOI.get( optimizer, MOI.RawOptimizerAttribute("abstract_controller") ) concrete_controller = MOI.get( optimizer, MOI.RawOptimizerAttribute("concrete_controller") )
@@
In Dionysos.jl, we can also generate visualizations thanks to the implemented recipes. In Code LABEL:lst:visu, we visualize the constructed abstract system, and it gives the plot in Figure 7. {lstlisting}[ language = Julia, numbers=left, label=lst:visu, caption=The recipe implemented for the abstraction is used to visualize the abstract system. ] fig = plot(aspect_ratio=:equal) title!(fig, "Abstraction"); plot!( fig, abstract_system; arrowsB = true, cost = false )
Figure 7: Result of Code LABEL:lst:visu.
@@ 6 Benchmarking
Table 2: Comparaison between SCOTS, CoSyMA and Dionysos.jl (AB.UniformGridAbstraction solver from Table 1) for the DC-DC converter example. Dionysos.jl outperforms SCOTS and CoSyMA with and without prior knowledge of the system’s incrementally stable property.
@@
Table 3: Comparaison between SCOTS, CoSyMA and Dionysos.jl (AB.UniformGridAbstraction solver from Table 1) for the path planning example. Dionysos.jl outperforms SCOTS.

### v2

@@ 1 Introduction
To address this issue, we introduce Dionysos.jl [^1] The package is open source and available on GitHub at: https://github.com/dionysos-dev/Dionysos.jl., a modular Julia [9] package for solving optimal control problems using state-of-the-art techniques, from control, optimization, and machine learning, for complex systems. It is built on top of different Julia packages such as JuMP.jl and MathOptInterface.jl, and features optimal control problem definitions and several abstraction-based methods to solve them. Dionysos.jl is the software of the ERC project Learning to control (L2C) [^2] European Reasearch Council (ERC) under the European Union’s Horizon 2020 research and innovation program under grant agreement No 864017 - L2C.. It implements new fundamental techniques that have been designed for smart abstraction [16, 30, 20, 6, 15] as part of the L2C project, with the goal of designing control techniques that would come with guarantees of safety and efficiency, while at the same time being able to take into account non-standard constraints or information, e.g. coming from first principles in physics, a logical specification translating some legal regulations, or some recommendation from a human being.
@@ 2.1 Control framework
A tuple $(\boldsymbol{x},\boldsymbol{u})\in\mathcal{X}^{T+1}\times\mathcal{U}^{T}$ is a trajectory of length $T$ of the system $\mathcal{S}=(\mathcal{X},\mathcal{U},F)$ starting at $x(0)$ if $T\in\mathbb{N}\cup\{\infty\}$, $x(0)\in\mathcal{X}$, $\forall k\in\{0,\ldots,T-1\}:u(k)\in\mathcal{U}(x(k))$ and $x(k+1)\in F(x(k),u(k))$. The set of trajectories of $\mathcal{S}$ is called the behavior of $\mathcal{S}$, denoted $\mathcal{B}(\mathcal{S})$.
@@
We now define the control problem. {defi} Consider a system $\mathcal{S}=(\mathcal{X},\mathcal{U},F)$. A specification $\Sigma$ for $\mathcal{S}$ is defined as any subset $\Sigma\subseteq(\mathcal{X}\times\mathcal{U})^{*}\cup(\mathcal{X}\times\mathcal{U})^{\infty}$. It is said that system $\mathcal{S}$ satisfies the specification $\Sigma$ if $\mathcal{B}(\mathcal{S})\subseteq\Sigma$. A system $\mathcal{S}$ together with a specification $\Sigma$ constitute a control problem $(\mathcal{S},\Sigma)$. Additionally, a controller $\mathcal{C}$ is said to solve the control problem $(\mathcal{S},\Sigma)$ if $\mathcal{C}\times\mathcal{S}$ satisfies the specification $\Sigma$.
@@ 2.3 Smart abstraction
Figure 3: Comparison of piecewise constant and state-dependent controllers. The red region illustrates $F_{1}(x_{1},u)$ for all $x_{1}\in R^{-1}(x_{2})$, given $u\in\mathcal{U}_{1}$ where $u\in\mathcal{U}_{2}(x_{2})$. The blue region shows $G_{1}(x_{1})=F_{1}(x_{1},\kappa(x_{1}))$ for all $x_{1}\in R^{-1}(x_{2})$, where $\kappa\in\mathcal{U}_{2}(x_{2})$ is a local state-dependent controller $\kappa:\mathcal{X}_{1}\rightarrow\mathcal{U}_{1}$. Left: The two-dimensional concrete system with its state space discretization. Right: The corresponding abstract system, highlighting the non-deterministic transition $F_{2}(x_{2},u)$ and the deterministic transition $F_{2}(x_{2},\kappa)$.
@@
which requires that all states in the initial set $\mathcal{X}_{I}$ remain (safely) in the set $\mathcal{X}_{S}$ forever. In addition to the specifications, a state cost function $c:\mathcal{X}\rightarrow\mathbb{R}_{+}$ evaluating the cost of being in a state $x$, and a transition cost function $t:\mathcal{X}\times\mathcal{U}\rightarrow\mathbb{R}_{+}$ quantifying the cost of transitioning from one state $x$ to another with control input $u$, can be supplied to the control problem. The objective is then to design a controller that satisfies the specification while minimizing the cumulative cost. Discretization templates: The quantizer $R$ can either be a (partial or not) partition or cover of the concrete state space. In addition, given a continuous state space $\mathcal{X}\subseteq\mathbb{R}^{n}$, Dionysos.jl supports two types of sets for the quantizer: hyperrectangles (Definition 3), and ellipsoids (Definition 3).
{defi}
A hyperrectangle of center $c\in\mathbb{R}^{n}$ and half-lengths $h\in\mathbb{R}^{n}_{+}$ is defined as
$\operatorname{H}(c,h)=\{x\in\mathbb{R}^{n}\mid|x_{i}-c_{i}|\leq h_{i}\text{ for }i=1,\ldots,n\}.$
{defi}
An ellipsoid with center $c\in\mathbb{R}^{n}$ and shape defined by $P\in\mathbb{S}_{+}^{n}$ is defined as
$\operatorname{E}(c,P)=\{x\in\mathbb{R}^{n}\mid(x-c)^{\top}P(x-c)\leq 1\}.$
@@ 4.1 The System module
The System module contains mathematical descriptions of controlled systems (see Definition 2.1), controllers (see Definition 2.1) and trajectories. It is an extension of MathematicalSystems.jl and HybridSystems.jl [29], and complements the latter with more specific system definitions.
@@ 4.2 The Problem module
Note that, in Dionysos.jl, the obstacles are encoded as part of the domain of the system [^3] A description of the Domain module can be found in the package documentation.. The second structure is SafetyProblem{S,XI,XS,T<:Real}, and implements an invariance problem. Its fields are
@@ 4.3 The Optim module
A simple example of an implementation of a classical abstraction-based method is given in the following. In Code LABEL:lst:optimizer_optimize!, the steps given in Figure 1 are followed. In Code LABEL:lst:optimizer_struct, the structure for the optimizer is defined. To be initialized, it needs a concrete problem to solve, as well as a discretization for the state space and the input space. {lstlisting}[ language = Julia, numbers=left, numbersep=6pt, xleftmargin=0.8em, label=lst:optimizer_struct, caption=Definition of an abstraction-based strategy structure implementing the abstract type MOI.AbstractOptimizer., captionpos=b ] mutable struct OptimizerT<:MOI.AbstractOptimizer concrete_problem::Union Nothing, PR.OptimalControlProblem, PR.SafetyProblem abstract_problem::Union Nothing, PR.OptimalControlProblem, PR.SafetyProblem abstract_controller::Any concrete_controller::Any state_grid::Any input_grid::Any function OptimizerT(cp,sg,ig) where T return newT( cp, nothing, nothing, nothing, nothing, sg, ig ) end end {lstlisting}[ language = Julia, numbers=left, numbersep=6pt, xleftmargin=0.8em, label=lst:optimizer_optimize!, caption=Definition of the MOI.optimize! function for the strategy Optimizer defined in Code LABEL:lst:optimizer_struct., captionpos=b ] function MOI.optimize!(opt::Optimizer) # Build the abstraction abstract_system = build_abs_system( opt.concrete_problem.system, opt.state_grid, opt.input_grid, ) # Build the abstract problem abstract_problem = build_abs_problem( opt.concrete_problem, abstract_system ) opt.abstract_problem = abstract_problem # Solve the abstract problem abstract_controller = solve_abs_problem( abstract_problem ) opt.abstract_controller = abstract_controller # Solve the concrete problem opt.concrete_controller = solve_conc_problem( abstract_system, abstract_controller ) end
@@ 5 Numerical example
Let $E(c,P)\subseteq\mathbb{R}^{n}$ be an ellipsoid with center $c\in\mathbb{R}^{n}$ and shape defined by $P\in\mathbb{R}^{n\times n}$ as defined in Definition 3.
@@
The system in Example 5, as well as other examples, are already defined in Dionysos.jl/problems. The OptimalControlProblem defining the problem in Example 5 can be found in the NonLinear module in problems/non_linear.jl. In Code LABEL:lst:ex_system, we import this problem. {lstlisting}[ language = Julia, numbers=left, numbersep=6pt, xleftmargin=0.8em, label=lst:ex_system, caption=The problem stated above is imported from the problems directory., captionpos=b ] concrete_problem = NonLinear.problem() concrete_system = concrete_problem.system
@@
We choose to use AB.LazyEllipsoidsAbstraction (see Table 1) to solve this problem, a smart abstraction method that constructs a memoryless concretization relation ($\operatorname{MCR}$) that partially covers the state space with ellipsoids. In Code LABEL:lst:ex_solver, we instantiate the corresponding Optimizer, then set all the needed fields, including concrete_problem, the reach-avoid problem stated above. For the sake of conciseness, we do not state the purpose of the other fields, and we gather them in the variable other_parameters in the code. {lstlisting}[ language = Julia, numbers=left, numbersep=6pt, xleftmargin=0.8em, label=lst:ex_solver, caption=The Optimizer is instantiated, and the solver parameters are set., captionpos=b ] optimizer = MOI.instantiate( AB.LazyEllipsoidsAbstraction.Optimizer ) AB.LazyEllipsoidsAbstraction.set_optimizer!( optimizer, concrete_problem, other_parameters… )
@@
In Code LABEL:lst:ex_solve, we solve the problem and retrieve the corresponding abstract system, abstract problem and concrete controller (see Figure 1). Note that every operation follows the MathOptInterface.jl syntax. {lstlisting}[ language = Julia, numbers=left, numbersep=6pt, xleftmargin=0.8em, label=lst:ex_solve, caption=The problem is solved, and the result is extracted from the solver., captionpos=b ] MOI.optimize!(optimizer) abstract_system = MOI.get( optimizer, MOI.RawOptimizerAttribute("abstract_system") ) abstract_problem = MOI.get( optimizer, MOI.RawOptimizerAttribute("abstract_problem") ) abstract_controller = MOI.get( optimizer, MOI.RawOptimizerAttribute("abstract_controller") ) concrete_controller = MOI.get( optimizer, MOI.RawOptimizerAttribute("concrete_controller") )
In Code LABEL:lst:ex_simulate, we simulate a concrete closed-loop trajectory starting from an initial point $x_{0}$ sampled from the initial set $\mathcal{X}_{I}$. {lstlisting}[ language = Julia, numbers=left, numbersep=6pt, xleftmargin=0.8em, label=lst:ex_simulate, caption=Simulate a closed-loop trajectory., captionpos=b, ] x0 = UT.sample(concrete_problem.initial_set) reached(x) = x ∈concrete_problem.target_set trajectory = ST.get_closed_loop_trajectory( concrete_system, concrete_controller, concrete_problem.transition_cost, x0; stopping = reached, )
@@
In Dionysos.jl, we can also generate visualizations thanks to the implemented recipes. In Code LABEL:lst:visu, we visualize the constructed abstract system, and it gives the plot in Figure 6. {lstlisting}[ language = Julia, numbers=left, numbersep=6pt, xleftmargin=0.8em, label=lst:visu, caption=The recipe implemented for the abstraction is used to visualize the abstract system. The cost is set to true to plot the upper-bound on the cost to reach the target., captionpos=b, ] fig = plot(aspect_ratio=:equal) plot!(abstract_system; arrowsB = true, cost = true) plot!(concrete_problem.target_set; color = :red) plot!(trajectory; color = :blue) plot!(concrete_problem.initial_set; color = :green)
Figure 6: Results of Code LABEL:lst:visu for different meta-parameters (other_parameters) of the solver AB.LazyEllipsoidsAbstraction. The initial set $\mathcal{X}_{I}$, target set $\mathcal{X}_{T}$, and obstacle set $\mathcal{X}_{O}$ are shown in green, red, and black, respectively. The value function, which provides an upper bound on the cost to reach the target, is shown in a blue color map, with the closed-loop system’s trajectory also displayed in blue.
@@ 6 Benchmarking
Table 2: Comparison between SCOTS, CoSyMA and Dionysos.jl (AB.UniformGridAbstraction solver from Table 1) for the DC-DC converter example. Dionysos.jl outperforms SCOTS and CoSyMA with and without prior knowledge of the system’s incrementally stable property.
@@
Table 3: Comparison between SCOTS, CoSyMA and Dionysos.jl (AB.UniformGridAbstraction solver from Table 1) for the path planning example. Dionysos.jl outperforms SCOTS.